# UAV-borne Mapping Algorithms for Canopy-Level and High-Speed Drone Applications

### v1

@@ 3.3.2 Geometric Accuracy
With the correspondences found using the ICP algorithm, the geometric accuracy of the reconstructed map is measured by the distance between corresponding points in the ground truth 3D model and the reconstructed 3D map. Root Mean Square Error (RMSE) is used as a metric of geometric accuracy. RMSE is a measure of the average error between the reconstructed and ground truth positions. It penalizes larger errors more heavily due to the squared term. The equation for calculating RMSE is:
$RMSE=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\mathbf{P}_{i}-\mathbf{T}_{i}\right)^{2}}$ (5)
where $N$ is the total number of points in the reconstructed point cloud, $\mathbf{P}_{i}$ is the position of the $i$-th reconstructed point, and $\mathbf{T}_{i}$ is the ground truth (reference) position of the corresponding $i$-th point.
@@
$\sigma=\sqrt{\frac{\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}{N}}$ (6)
where $\bar{x}$ is the mean of the errors.
@@ 4 Results
Simulated experiments were conducted using a virtual model of the UNC Charlotte campus near the football stadium. The virtual model of this environment was generated using the Unreal Engine and the Cesium plug-in. The imaging setup on the simulated multirotor comprised a pair of stereo cameras. For DSO (SfM), only data from the left camera was utilized, while DSOL and SDSO (stereo reconstruction) utilized the stereo data. Additionally, a depth sensor on the multirotor provided ground truth depth information and scene point cloud. The multirotor maintained an average speed of 16.5 m/s, with a maximum horizontal speed of 20 m/s during the flight. Fig. 7(a) illustrates the simulated drone’s flight over the UNC Charlotte football stadium, covering a 3-minute flight duration and capturing 1027 RGB-D frames. Sample images from the drone’s simulated RGB and depth sensors are displayed in Figs. 7(b) and 7(c).
(a) A multirotor flying over a simulated virtual model of UNC Charlotte campus.
(b) A RGB image captured by the drone camera.
(c) A depth image captured by the drone camera.
Figure 8: A simulated UNC Charlotte campus world.
The experiments were conducted on an Intel i7-12700KF CPU. The official implementation of DSO and DSOL were used for evaluation. An open-source implementation of SDSO [25] based on the official DSO with stereo integration was employed, as no official implementation is available. All implementations adhered to the original configuration with the default parameters.
@@ 4.1 Mapping Accuracy Evaluation
The ground truth point cloud of the experimental scene is shown in Fig. 9. This was generated by transforming the point cloud of each RGB-D frame to the location of the multirotor obtained from the ground truth odometry provided by AirSim. This point cloud was used as the ground truth geometry to evaluate the mapping accuracy of different algorithms.
@@
Fig. 10 shows the point cloud respectively reconstructed by DSO, SDSO, and DSOL. All three maps reflect the experimental scene. However, there are a lot richer structures in both DSO and DSOL maps than in the DSOL map. Detailed analysis is provided in the following sections.
(a) DSO
(b) SDSO
(c) DSOL
Figure 10: Point clouds generated by DSO (left), SDSO (middle), and DSOL (right). DSO and SDSO generated much more point clouds than DSOL. The color of the points is represented by the grayscale color of the scene point.
@@ 4.1.1 Quantitative Analysis
GT DSO SDSO DSOL
points 236032175 204345 212179 6662
correspondences - 172862 175567 2670
map scale - 35.98 1.01 1.11
RMSE - 0.155 0.164 0.229
std - 0.110 0.114 0.141
Table 1: Mapping accuracy of three mapping algorithms. DSO and SDSO maps contain $\sim$30$\times$ more points than the DSOL map. The map scales are obtained by estimating the scale factor from the reconstructed point cloud to the ground truth. DSO, as a monocular mapping algorithm, is not able to correctly estimate the scale of the map.
Table 1 details the density and accuracy characteristics of mapping results obtained from different algorithms. In our experimental setup, we utilized the Iterative Closest Point (ICP) algorithm to align the reconstructed point clouds (source) with the ground truth point cloud (target). The specified search radius, representing the distance from each point in the source point cloud within which the ICP algorithm seeks a corresponding point in the target point cloud, is set at 0.5 meters. Once all correspondences are established, the scale factor from the source to the target point cloud, along with the fitting error, can be derived. The “GT” column provides information about the ground truth point cloud, while the three subsequent columns present statistics related to the reconstructed point clouds. Notably, DSO and SDSO maps encompass approximately 30 times more points than the DSOL map, aligning with the observations in Fig. 10. The contrast in the correspondence sets is more pronounced, with DSO and SDSO revealing approximately 65 times more correspondences than DSOL. The “map scale” reflects the scale factor from the reconstructed point cloud to the ground truth, with DSO unable to accurately estimate scale as a monocular mapping algorithm. The increased number of correspondences likely provides a more robust basis for establishing the scale between the reconstructed point cloud and the ground truth, resulting in a more accurate map scale for SDSO. The “RMSE” and “std” rows, namely the Root Mean Square Error and standard deviation, reveal noteworthy distinctions in accuracy and consistency. DSO exhibits the lowest RMSE (0.155) and a small std (0.110), indicating both high accuracy and consistency. SDSO follows closely with an RMSE of 0.164 (5.81% higher) and a comparable std of 0.114 (3.64% higher). In contrast, DSOL demonstrates a higher RMSE of 0.229 (47.74% higher than DSO and 39.63% than SDSO) and a larger std of 0.141 (28.18% higher than DSO and 23.68% than SDSO), suggesting lower accuracy and greater variability. While DSO and SDSO prove more accurate and consistent in their mapping results, DSOL lags in terms of both precision and reliability.
[…]
Figure 11: The histogram of the fitting error. The error is measured as the distance from the points reconstructed by mapping algorithms to their corresponding points in the ground truth point cloud.
[…]
@@
Figure 12: Distribution of point depth in the reconstructed maps (left) and reconstruction RMSE of points at different depth ranges (right).
@@ 4.1.2 Qualitative Analysis
A qualitative examination of the reconstructed maps from DSO, SDSO, and DSOL, as illustrated in 13, unveils notable distinctions in their alignment with the ground truth. DSO and SDSO, with their significantly higher point densities, exhibit commendable accuracy in capturing the details of the road network, showcasing intricate and well-aligned representations. The enhanced point density, particularly evident in the football stadium region, allows for a more detailed reconstruction, resulting in superior alignment with the ground truth in this specific area. Conversely, DSOL, characterized by a sparser point cloud, displays a more conservative reconstruction, which translates into a reduced level of detail, particularly in complex structures like the football stadium. Although DSOL shows commendable alignment on the roads, its limitations in point density become apparent in areas with more intricate features.
(a) DSO
(b) SDSO
(c) DSOL
Figure 13: Reconstructed point clouds (blue) overlaid with the ground truth point cloud (actual color): DSO (left), SDSO (middle), and DSOL (right). All three reconstructed point clouds have been scaled by the factors in Table. 1.
@@ 4.2 Computational Cost Evaluation
This section is dedicated to a comprehensive examination of the computational cost exhibited by the three mapping algorithms. The effectiveness of these algorithms is scrutinized with a particular focus on two critical aspects of mapping computational cost: keyframe creation time and frame tracking time. These metrics serve as crucial benchmarks in assessing the algorithms’ ability to swiftly and accurately generate keyframes, as well as tracking real-time camera pose changes during the mapping process. Through this meticulous examination, we seek to offer valuable insights that contribute to the informed selection and deployment of mapping solutions across a spectrum of real-world applications.
@@ 4.2.1 Keyframe Creation Time
[…]
The “total kfs” column in Table 2 provides insights into the aggregate number of keyframes generated by the three algorithms. DSOL, employing a simplified keyframe creation criterion, notably generates approximately $\sim$80fewer keyframes compared to its counterparts. Noteworthy is the observation that DSO creates 29 fewer keyframes than SDSO. This discrepancy is attributed to a delayed initialization of the DSO system, commencing at ${}^{\text{th}}$ frame in our experiments, in contrast to the immediate initialization of SDSO and DSOL. Such delay is also illustrated in Fig. 14 as the DSO curve starts later than SDSO and DSOL. DSO system’s initialization process relies on assigning random depth values to candidate points and predicting the initial camera movement pattern, demanding precise assumptions about initial depth values and camera motion. In contrast, mapping systems employing stereo cameras, such as SDSO and DSOL, leverage stereo matching for enhanced depth initialization, leading to increased accuracy. The divergence in keyframe quantities among the algorithms also mirrors the disparities in point cloud density depicted in Fig. 10, given that these points are derived from the keyframes.
total kfs min ($ms$) max ($ms$) mean ($ms$) std
@@
Table 2: Statistics of the keyframe creation results of three mapping algorithms. The “total kfs” column shows the total amount of the keyframes created by three algorithms. The “min”, “max”, and “mean” respectively show the minimum, maximum, and average time for keyframe creation. The “std” column denotes the standard deviation of the keyframe creation time.
@@ 4.2.2 Frame Tracking Time
[…]
Figure 15: Frame tracking time of different algorithms. The left figure shows the tracking time for all frames including keyframes and non-keyframes. The right figure provides a close-up of the DSO and SDSO tracking time for frames 290$\sim$310. The data is plotted as scatter points for a clear visualization.
5 Conclusions
This study delved into the context of canopy-level and high-speed drone applications. A thorough examination of various sensors underscored their unique strengths and challenges, guiding the selection of suitable devices for specific operational scenarios. The experiments centered on evaluating three prominent mapping algorithms—DSO, SDSO, and DSOL—in a simulated environment, providing valuable insights into the performance of these mapping algorithms. Each algorithm exhibits unique strengths and trade-offs, catering to specific requirements in UAV-based mapping scenarios. DSO, operating as a monocular mapping algorithm, demonstrates versatility in capturing scenes with a single camera, albeit with limitations in scale estimation. SDSO, incorporating stereo depth perception, excels in accuracy and spatial fidelity, as evidenced by its superior point cloud density and detailed reconstructions, particularly in complex structures like the football stadium. On the other hand, DSOL, designed for efficiency, streamlines the mapping process, offering reliable reconstructions with reduced computational demands. The findings emphasize the importance of algorithm selection based on the specific needs of UAV-borne mapping applications, considering factors such as scale estimation, spatial accuracy, and computational efficiency. Future work may involve refining these algorithms for optimized performance in diverse environments, ultimately contributing to advancements in UAV-based mapping for canopy-level and high-speed drone applications. This study contributes to the ongoing discourse on mapping algorithms, providing valuable insights for researchers and practitioners navigating the dynamic landscape of UAV applications in remote sensing and environmental monitoring.

### v2

@@ 3.3.2 Geometric Accuracy
With the correspondences found using the ICP algorithm, the geometric accuracy of the reconstructed map is measured by the distance between corresponding points in the ground truth 3D model and the reconstructed 3D map. The mean of this distance of all the corresponding points is used to evaluate the accuracy performance, calculated as follows:
$\bar{x}=\frac{1}{N}\sum_{i=1}^{N}||\mathbf{P}_{i}-\mathbf{T}_{i}||$ (5)
where $N$ is the total number of corresponding points, $\mathbf{P}_{i}$ is the position of the $i$-th corresponding point, and $\mathbf{T}_{i}$ is the ground truth (reference) position for that point.
@@
$\sigma=\sqrt{\frac{\sum_{i=1}^{N}(||\mathbf{P}_{i}-\mathbf{T}_{i}||-\bar{x})^{2}}{N}}$ (6)
3.3.3 Computational Cost
To assess the computational efficiency of mapping algorithms, two key metrics were focused on in this article: keyframe creation time and frame tracking time. These metrics were selected to provide insights into the mapping speed of the algorithms.
Keyframe Creation Time: Keyframe creation time quantifies the time required to identify keyframes during the mapping process. Keyframe creation is arguably the most time-consuming process of the mapping pipeline, often 5–10$\times$ slower than tracking [qu2022dsol]. Creating too many keyframes will cause the system to eventually lag behind the frame rate. Keyframe creation time reflects the computational efficiency of map reconstruction.
Frame Tracking Time: Frame tracking time represents the duration required for the algorithms to process and track individual frames with respect to the keyframes. This metric reflects the algorithm’s ability to track and update the mapping information in real-time.
These two metrics collectively provide a comprehensive evaluation of the computational cost of mapping algorithms. The results of these evaluations are discussed in Section 4.2, providing the relative efficiency and performance trade-offs among the implemented algorithms.
@@ 4 Results
The experimental scene was a virtual model of the UNC Charlotte campus near the football stadium. The model was generated using AirSim and the underlying Unreal Engine in combination with the Cesium Tiles plugin.
Experiments were conducted on an Intel i7-12700KF CPU. The implementations of DSO and DSOL that were made available on GitHub by the authors were used [DSOimplementation, DSOLimplementation]. SDSO implementation by the authors is not available so an open-source third-party implementation on GitHub was chosen [SDSOimplementation]. All implementations adhered to the original configuration optimized by their authors for accuracy and/or speed performance including the number of active keyframes and maximal tracking points per frame. Customized modifications made to all three algorithms respectively for collecting experimental data include the following: (1) saving the generated point cloud to a PCD file; (2) saving the keyframe ID and associated creation time to a text file; (3) saving the frame ID and associated tracking time to a text file.
Experiments consist of a simulated quadrotor vehicle that traverses the virtual scene at heights ranging from 12 m to 20 m and at speeds ranging from 16.5 m/s to 20 m/s. During the flight camera sensor telemetry was recorded from a stereo pair of camera mounts to the UAV chassis. Algorithms processed the telemetry to generate mapping data for the environment. The SfM algorithm (DSO) used data from the left camera of the stereo rig while DSOL and SDSO (stereo reconstruction) utilized all the available image data. The left camera is chosen to define the sensor coordinate system and the noiseless depth sensor is co-located with the left camera to record ground truth depth for each pixel measured within the view of the left camera. Figure 10a illustrates the simulated drone’s flight over the UNC Charlotte football stadium, covering a 3-minute flight duration and capturing 1027 RGB stereo pair frames and associate ground truth depth. Sample images from the drone’s simulated RGB and depth sensors are displayed in Figure 10b,c.
a
b
c
Figure 10: A simulated UNC Charlotte campus world. (a) A quadrotor flying in a virtual model of UNC Charlotte; (b) A RGB image captured by the drone camera; (c) A depth image captured by the drone camera .
@@ 4.1 Mapping Accuracy Evaluation
Using the methods of Section 3.2.3, a ground truth point cloud of the experimental scene was calculated which is shown in Figure 11. This point cloud serves as the ground truth geometry to evaluate the mapping accuracy of different algorithms. Figure 12 shows the point cloud respectively reconstructed by DSO, SDSO, and DSOL. All three maps qualitatively encode the shape and size of objects from the experimental scene. However, the point density and scene details of DSO and DSOL maps outperform the DSOL map.
@@
a
b
c
Figure 12: Point clouds generated by (a) DSO , (b) SDSO, and (c) DSOL. DSO and SDSO generated much more point clouds than DSOL. The color of the points is represented by the grayscale color of the scene point.
@@ 4.1.1 Quantitative Analysis
Table 1 details the density and accuracy characteristics of mapping results obtained from different algorithms. The ICP algorithm was used to align estimate maps with the ground truth point cloud. Criteria for alignment convergence and correspondence calculation included the following: (1) a search radius of 0.5 m, (2) algorithm termination criteria which are triggered when either the RMSE of corresponding points changes by less than 0.00001 or the maximum number of iterations exceeds 1500. Alignment results allow for statistics to be tabulated on the point cloud accuracy for each algorithm. DSO, as a monocular SfM algorithm, is unable to estimate the scene scale accurately. The scale factor was estimated from the ICP algorithm for DSO and the DSO mapping result (point cloud) was scaled by the estimated factor for accuracy evaluation. For SDSO and DSOL, the scale factor was not estimated since the scale of the scene can be derived from the stereo data. In our experiments, the estimated map scale of DSO was 35.98. Table 1 captures key statistics of the alignment process for the three algorithms evaluated. Each row of this table is explained below:
  • points: Total points in the reconstructed point cloud.
  • correspondences: Total amount of correspondences.
  • mean: Mean of the distance between all corresponding points.
  • std: Standard deviation of the distance between all corresponding points.
Table 1: Quantitativeevaluation of the point clouds generated by three mapping algorithms. DSO and SDSO maps contain more points and correspondences than the DSOL map. The error statistics (“mean” and “std”) indicate both higher accuracy and consistency in the DSO and SDSO mapping results than DSOL. The DSO map was scaled by a factor (35.98) estimated from the ICP algorithm.
DSO SDSO DSOL
points 204345 212179 6662
correspondences 172862 183460 2799
mean (m) 0.110 0.110 0.177
std (m) 0.110 0.111 0.145
Notably, DSO and SDSO maps, similar to each other in the total amount of points, encompass $\sim$30 times more points than the DSOL map. This aligns with the observations in Figure 12. The difference in the correspondence sets is more pronounced, with DSO and SDSO revealing $\sim$65 times more correspondences than DSOL. DSO, after scaling, follows very closely to SDSO in terms of mapping accuracy performance, while both of them exhibit a 60.9% lower mean error than DSOL and a 31.8% smaller standard deviation. DSO and SDSO prove more accurate and consistent in their mapping results, while DSOL lags in terms of both precision and reliability.
Figure 13 depicts the distribution of the distance between corresponding map locations for the three algorithms. DSO and SDSO exhibit similar distributions and most 3D measurements lie within 0.15 m to their corresponding location in the ground truth model. In contrast, DSOL has significantly fewer points within the 0.15 m distance range and a nearly constant number of points having similar errors for greater distances. This supports the mapping accuracy results shown in Table 1.
Figure 14 indicates the capability of each algorithm to estimate large depths from a given viewpoint (Figure 14a) and expected error for a depth estimate for each depth (Figure 14b) where depths have been binned to 5 m intervals for tabulation.
Figure 14a shows the distribution of depth values for the keyframes of the trajectory which are responsible for generating depth values. Figure 14a indicates that a majority of depth estimates range from 20 m to 60 m. One can also see that DSO is capable of generating estimates at larger depths than the two other algorithms (see ranges 100–130 m). DSOL tends to reconstruct points within 60 m and shows a slightly bimodal behavior with a high population of measurements in the 60–100 m range which may be an artifact due to the experimental context. Figure 14b portrays the expected depth error in each keyframe. Figure 14a also lacks any presence of short ranges. This can be attributed to flying at low-altitude where most data is further than 10 m away.
Figure 13: Quantitative analysis of the distribution of the closest point distances for correspondences from the mapping results to the ground truth point cloud. DSO and SDSO have similar distributions while the mean correspondence error is larger in DSOL.
a
b
@@
Figure 14: (a) Distribution of point depth in the reconstructed maps and (b) the average distance between the matched points at different depth ranges.
Figure 14b, shows the expected depth error for estimate depths. Inspection of the results for distances of 20–40 m, the reconstruction error of DSOL is approximately 0.15 m per point while DSO and SDSO are close to each other having an error of approximately 0.085 m. DSO outperforms SDSO across most depth ranges with slightly smaller distance measurements. Additionally, the error distributions exhibit a quadratic growth pattern as predicted by theoretical models as described in Figure 2. High error is noted at short ranges of less than 25 m. This can be attributed to a lack of sufficient supporting image data due to the high velocity of the UAV. Surfaces close to the vehicle move quickly through the field of view and exhibit more motion artifacts leading to higher depth estimation error.
Figure 14 indicates that DSO exhibits lower error values across all ranges yet has fewer points. This can be attributed to a strong filter on the acceptable point depth covariance for map points within the algorithm. SDSO and DSOL exhibit lower accuracy compared to those produced by DSO.
@@ 4.1.2 Qualitative Analysis
Figure 15 shows reconstructed maps from DSO, SDSO, and DSOL. A qualitative examination of these results unveils notable distinctions in their alignment with the ground truth. DSO and SDSO, with their significantly higher point densities appear to exhibit good accuracy as evidenced by the details of the road network that have been captured and include intricate and well-aligned geometries, e.g., road curbs. The enhanced point density, particularly evident in the football stadium region, allows for a more detailed reconstruction and appears to provide better alignment results relative to ground truth here. Conversely, DSOL, characterized by a sparser point cloud provides a reduced level of detail, particularly in complex structures like the football stadium. Although DSOL shows good alignment for roads, the sparsity of the estimate limits the map details.
a
b
c
Figure 15: Reconstructed point clouds (blue) overlaid with the ground truth point cloud (actual color): (a) DSO , (b) SDSO, and  (c) DSOL. DSO point cloud has been scaled by the factor estimated by ICP.
@@ 4.2 Computational Cost Evaluation
Computation cost for the considered algorithms considers the resources required by two critical mapping algorithm functions cost: (1) keyframe creation time and (2) frame tracking time. These metrics serve as crucial benchmarks in assessing the algorithms’ ability to swiftly and accurately generate keyframes, as well as tracking real-time camera pose changes during the mapping process. Through this examination, we seek to offer valuable insights that contribute to the informed selection and deployment of mapping solutions for low-altitude UAV flights, particularly for high-speed applications.
@@ 4.2.1 Keyframe Creation Time
Figure 16 illustrates the keyframe creation time for DSO, SDSO, and DSOL. It can be seen that SDSO requires the most time to create a keyframe, averaging $\sim$220.73 ms per keyframe, as reported in Table 2. DSO incurs lower computational cost for keyframe creation since the stereo disparity map estimation algorithm is not required resulting in keyframe times averaging around $\sim$200.28 ms per keyframe. DSOL requires $\sim$7.39 ms per keyframe which is approximately 30 times faster than competing approaches. This can be attributed to the simplified keyframe creation process facilitated as a combination of a simplified disparity computation algorithm and parallel processing. The columns in Table 2 delineate the statistical distribution of keyframe creation times, including the minimum, maximum, and mean values, with the “std” column denoting the standard deviation. In summary, SDSO necessitates 10.21% more keyframe creation time than DSO and 2886.87% more than DSOL, while DSO requires 2610.15% more time than DSOL.
Table 2: Statistics of the keyframe creation results of three mapping algorithms. The “total kfs” column shows the total amount of the keyframes created by three algorithms. The “min”, “max”, and “mean”, respectively, show the minimum, maximum, and average time for keyframe creation. The “std” column denotes the standard deviation of the keyframe creation time.
Total kfs min (ms) max (ms) mean (ms) std (ms)
@@
Table 2 contains data that provides quantitative measures for the aggregate number of keyframes generated by the three algorithms (“total kfs” column). DSOL generates approximately $\sim$80fewer keyframes compared to its counterparts which can be attributed to slightly more restrictive requirements for keyframe creation. Noteworthy is the observation that DSO creates 29 fewer keyframes than SDSO. This discrepancy is attributed to a delayed initialization of the DSO system, commencing at the 50th frame in our experiments, in contrast to the immediate initialization of SDSO and DSOL. Such delay is also illustrated in Figure 16 as the DSO curve starts later than SDSO and DSOL. The DSO initialization process relies on assigning random depth values to candidate points and predicting the initial camera movement pattern, demanding precise assumptions about initial depth values and camera motion. In contrast, mapping systems employing stereo cameras, such as SDSO and DSOL, leverage stereo matching for enhanced depth initialization, leading to increased accuracy. Divergence in keyframe quantities among the algorithms also mirrors the disparities in point cloud density depicted in Figure 12, given that these points are derived from the keyframes.
@@ 4.2.2 Frame Tracking Time
Figure 17 depicts the frame tracking time across various algorithms, employing scatter points for visualization. Results show the very high performance achieved by DSOL which requires very little computation for each tracked frame. In the case of DSO and SDSO, the tracking time is stratified into two distinct regions. The upper region, requiring approximately $\sim$60 ms for tracking, corresponds to keyframes, while the lower region, with an average tracking time of $\sim$20 ms per frame, pertains to non-keyframes. This 3×3$\times$3 × difference in tracking time arises from the creation of a new keyframe, where existing point tracks must be terminated and a collection of new point tracks must be initialized incurring significant computational cost to transfer the tracking information. Subsequent frames are then exclusively tracked to this keyframe, employing traditional two-frame direct image alignment methods. This stratification in tracking time offers insights into the computational demands associated with keyframe and non-keyframe tracking, highlighting the intricacies involved in SfM methods that must maintain accurate and efficient tracking across consecutive frames.
a
b
Figure 17: Frame tracking time of different algorithms. (a) shows the tracking time for all frames including keyframes and non-keyframes. The data are plotted as scatter points for a clear visualization. (b) shows the DSO and SDSO tracking time for frames 290$\sim$310 which corresponds to the region highlighted by the red box in (a).
Figure 17 also shows two apparent bands in the lower region for results of DSO and SDSO. The higher band characterizes the tracking time for frames immediately succeeding keyframes, while the lower band denotes the tracking time for other frames. A repeated pattern exists where $\sim$5 ms of addition time is required to process frames following keyframes. Figure 17b zooms into a subsection of the data associated with frame indices 290–310. Close examination of this phenomenon indicates that newly formed tracks require more time as the points of the initial keyframe have to be sorted into reliable and unreliable tracks thereby necessitating slightly more computation.
5 Conclusions
This paper presents a study on low-altitude and high-speed drone applications. An examination of various sensors underscored their strengths and challenges, guiding the selection of suitable devices for specific operational scenarios. The experiments centered on evaluating three prominent mapping algorithms—DSO, SDSO, and DSOL—in a simulated environment, providing valuable insights into the performance of these mapping algorithms. Each algorithm exhibits unique strengths and trade-offs, catering to specific requirements in UAV-based mapping scenarios. DSO, operating as a monocular mapping algorithm, demonstrates versatility in capturing scenes with a single camera, albeit with limitations in scale estimation. SDSO, incorporating stereo depth perception, excels in accuracy and spatial fidelity, as evidenced by its superior point cloud density and detailed reconstructions, particularly in complex structures like the football stadium. On the other hand, DSOL, designed for efficiency, streamlines the mapping process, offering reliable reconstructions with reduced computational demands. The findings suggest that, in cases where UAVs have limited computing resources, DSOL emerges as the optimal choice. For systems equipped with payload capacity and moderate compute resources, SDSO proves to be the most suitable option. When dealing with a single camera, DSO is the preferred choice for applications demanding dense mapping results.
Future work may involve refining these algorithms for optimized performance in diverse environments, ultimately contributing to advancements in UAV-based mapping for low-altitude and high-speed drone applications. This study contributes to the ongoing discourse on mapping algorithms, providing valuable insights for researchers and practitioners navigating the dynamic landscape of UAV applications in remote sensing and environmental monitoring.